# Triple Higgs boson production and electroweak phase transition in the two-real-singlet model

### v1

@@ 1 Introduction
“Electroweak baryogenesis” Kuzmin:1985mm ; Shaposhnikov:1986jp ; Shaposhnikov:1987tw ; Cohen:1990py ; Cohen:1993nk is a promising scenario that can generate this asymmetry. It demands the electroweak phase transition to be first order, providing the necessary departure from thermodynamic equilibrium (one of the Sakharov conditions Sakharov:1967dj ). Within the SM, the occurrence of a first-order phase transition (FOPT) requires the mass of the Higgs boson to lie below $\sim 70$ GeV Bochkarev:1987wf ; Kajantie:1995kf ; Kajantie:1996qd ; Csikor:1998eu , which is evidently inconsistent with the experimental observation of the Higgs boson mass of $\sim 125$ GeV. Extended scalar sectors can accommodate a FOPT that is, moreover, sufficiently strong to preserve any baryon asymmetry produced. We note further that the gravitational wave signal from a FOPT at the electroweak scale is in the right frequency range to be potentially observable with the LISA gravitational wave observatory Grojean:2006bp ; Caprini:2018mtu ; Caprini:2019egz ; LISACosmologyWorkingGroup:2022jok .
The SM scalar sector can be extended by an arbitrarily number of scalars Chung:2012vg ; Tenkanen:2022tly ; Zuk:2022qwx , but, arguably, the most minimal options are to add a complex or real singlet scalar field Espinosa:1993bs ; Benson:1993qx ; Vergara:1996ub ; Espinosa:2008kw ; Barger:2007im ; Espinosa:2011ax ; Cline:2012hg ; Curtin:2014jma ; Katz:2014bha ; Kanemura:2015fra ; Kurup:2017dzf ; Papaefstathiou:2020iag ; Papaefstathiou:2021glr , two real singlet fields Robens:2019kga ; Papaefstathiou:2020lyp ; Robens:2021zvr ; Robens:2022lkn ; Robens:2022hvu ; Robens:2023pax ; Robens:2023oyz , or additional doublet fields Land:1992sm ; Cline:1996mga ; Turok:1991uc ; Dorsch:2013wja ; Dorsch:2014qja ; Basler:2016obg ; Andersen:2017ika ; Dorsch:2017nza ; Bernon:2017jgv . The additional fields could induce a rich set of phenomena at colliders. In particular, if they mix with the SM-like Higgs boson, their corresponding physical states could be produced singly and decay into fermions or gauge bosons, yielding a signal in multi-lepton or jet searches for heavy resonances.
@@
Triple Higgs boson production is the first attainable process that allows to directly probe the quartic Higgs self-coupling in the scalar potential. Its observation would provide a clear signal for an extended scalar sector, as well as hints on its intricate structure, and possible rôle in electroweak baryogenesis. In fact, one might expect that there exists a correlation between enhanced multi-Higgs boson production and a strong FOPT necessary for baryogenesis, as they both require sufficiently large scalar couplings. Indeed, a FOPT can only arise if the changes to the SM scalar potential are significant, and previous studies of the electroweak phase transition in, e.g. the xSM, shows a clear preference for large couplings Kotwal:2016tex ; Alves:2018oct ; Papaefstathiou:2021glr . This intriguing observation constitutes the main motivation for the present article.
@@
We then study the possibility of a FOPT both in (1) the scenario with the necessary thermal barrier induced by the leading-order thermal mass corrections, as well as (2) a transition with a radiatively-generated barrier. Surprisingly, we find that the parameter space for large triple Higgs production boson and for a FOPT do not overlap. This result is intricately linked to the $\mathbb{Z}_{2}$-symmetric nature of the set-up and the choice of the vacuum today. Doubly-resonant triple Higgs boson production requires both additional scalars to mix with the Higgs field, and thus both to have a non-zero vacuum expectation value (vev) at zero temperature. If we give up on this requirement, and allow for one or both singlets to have zero vev, a FOPT becomes possible. Like in the xSM, di-Higgs boson production enhancement is possible if only one added scalar has a non-zero vev at low temperature. Unlike the xSM however, the TRSM introduces the possibility of a FOPT in this case.
The present article is organized as follows: in section 2 we discuss the scalar sector model of the $\mathbb{Z}_{2}$-symmetric two real singlet extension of the SM, discuss the conditions for large triple Higgs boson production, and their implications for the perturbativety bound. Section 3 contains our exploration of the parameter space of the model, including a list of updated benchmark points relevant to triple Higgs boson production. In section 4, we review the thermal history of the model, and analyze the possibilities for a FOPT both in the single- and multi-step scenario. We conclude in section 5. Several useful relations for the model are deferred to appendix A, such as the constraints on the boundedness of the potential, the connection between the mass and flavor basis, and the renormalization group equations (RGEs).
@@ 2.1 The two real singlet model
We point out that if one of the singlets has zero vev $v_{i}=0$ for $i=2,3$, the corresponding field $\phi_{i}$ does not mix with the other fields. This scenario does not offer novel collider phenomenology compared to the xSM. However, as we will see in section 4, it does have an interesting effect on cosmology, as it allows for a FOPT.
@@ 3 Updated scan for benchmark points
Figure 2: Scatter plot of the values of $M_{2}$ and $M_{3}$ for the 140 points with triple Higgs boson production cross section over 100 times the SM value.
@@ 4 Thermal history of the TRSM
It is well-known that loop corrections from particles in the plasma at finite temperature change the shape of the scalar field potential. As a result, the zero-temperature vacuum, in which all three scalar fields have a vev in our model, is not the vacuum state in the early universe. To end up in today’s vacuum, one or more PT(s) had to occur. These PT(s) can be of different types. In a FOPT the vev of one or several fields makes a discontinuous jump, due to a barrier in the potential. Instead, when the fields move continuously from the high-temperature phase to the zero-temperature phase, this is a second-order or cross-over PT. The electroweak PT in the SM is a cross-over Kajantie:1995kf ; Kajantie:1996mn ; Kajantie:1996qd ; Gurtler:1997hr ; Csikor:1998eu ; Aoki:1999fi , but in many BSM models the transition can be of a first-order type, see e.g. Caprini:2019egz and references therein for examples.
A FOPT is phenomenologically very interesting for two reasons: it can source a gravitational wave signal possibly observable by LISA, and it can provide the out-of-equilibrium dynamics necessary for electroweak baryogenesis. Both require the FOPT to be strong; for baryogenesis in particular, the SM Higgs vev should obtain a large value in the transition and
@@
It would be interesting to correlate the parameter space of enhanced triple Higgs production in the TRSM (and enhanced collider phenomenology in general) with the parameters for which a (sufficiently strong) FOPT can occur. For this, we study the effective potential at finite temperature using the framework of thermal field theory. As has become clear in recent years, accurate quantitative statements about PTs require careful many-loop computations Croon:2020cgk ; Niemi:2021qvp ; Gould:2023jbz ; Kierkla:2023von . However, to determine if a FOPT can occur in the TRSM, as a first approximation we can rely on much simpler expressions for the finite-temperature potential, that can be obtained from standard references, such as Dolan:1973qd ; Weinberg:1974hy , or textbooks Kapusta:2006pm ; Bellac:2011kqa .
@@ 4.1 The finite temperature effective potential
$V_{T,{\rm 1-loop}}(\phi,T)=\frac{T^{4}}{2\pi^{2}}\left[\sum_{\alpha=\Phi_{i},W,Z}n_{\alpha}J_{B}[m_{\alpha}^{2}(\phi_{i})/T^{2}]+n_{t}J_{F}[m_{t}^{2}(\phi_{i})/T^{2}]\right].$ (23)
@@ 4.2 Leading order phase transitions
Note that for the updated benchmark points discussed in Sec. 3, the global minimum does not always lie at the origin at high temperatures in the perturbative regime. Firstly, the scale where the masses become positive may lie above the perturbativity cut-off. Secondly, depending on the couplings, some of the thermal masses in eq. 26 can be negative, leading to a period of ‘symmetry non-restoration’ Weinberg:1974hy ; Mohapatra:1979vr ; Dvali:1995cc ; Meade:2018saz ; Baldes:2018nel ; Matsedonskyi:2020mlz . In the TRSM at least the Higgs thermal mass is always positive, because of the large top quark Yukawa contribution and the boundedness of the negative couplings eq. 54, and thus the electroweak symmetry always gets restored at high temperature. For concreteness, we will consider the case where all thermal masses are positive at temperatures below the perturbativity cutoff, and track the field dynamics as the universe cools down and all fields obtain a vev. If one or multiple thermal masses are already negative at the perturbativity scale, the analysis still applies, but starting from the point where there are already some non-zero vevs.
The PT dynamics in the TRSM mirrors that of the xSM, analyzed in e.g. Espinosa:2007qk ; Profumo:2007wc ; Cline:2009sn ; Espinosa:2011ax ; Curtin:2014jma ; Alves:2018jsw ; Gould:2019qek . In the well-studied two-step scenario Land:1992sm , only the Higgs field has a non-zero vev in the vacuum $(h,s)=(v_{h},0)$. Depending on the parameters, it is possible to have a two-step transition, where the first transition $(0,0)\to(0,s)$ is continuous, but the 2nd step $(0,s)\to(h,0)$ is a FOPT. However, if both fields have a vev in the zero-temperature vacuum $(h,s)=(v_{h},v_{s})$, all possible transitions are second order/cross over as they are always from a saddle point or maximum to a minimum. This has also been demonstrated in Espinosa:2011ax ; Ghorbani:2020xqv .
@@ 4.2.2 Phase transition dynamics
As the temperature drops further, $m_{1}^{2}$ becomes more negative and the extrema separate continuously from each other. The transition from $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$ is thus a continuous PT from a saddle point to a minimum, not a FOPT.
Finally, we note that there can never be two axial minima at the same time. For $m_{i}^{2},m_{j}^{2}<0$ the $\mathbf{x}_{i},\mathbf{x}_{j}$ extrema both exist. A necessary condition for both axial extrema to be a minimum is that $h_{jj}(\mathbf{x}_{i})>0$ and $h_{ii}(\mathbf{x}_{j})>0$, which can only be satisfied for $c_{ij}>0$, which then implies $C_{ij}<0$, which is at odds with our choice of vacuum eq. 38. Therefore, there cannot be a FOPT between different axial minima.
@@
We have already seen that two axial, two planar, and an axial and planar extrema can never be minima at the same time and that the origin is not a minimum once the other extrema emerge. Now our final analysis of the bulk minimum reveals that once it comes into existence, the planar (but also all other) extrema are not minima. Hence, as depicted in the left panel of fig. 3, all possible transitions are smooth and continuous, and no FOPT arises. This conclusion on the absence of a FOPT is based on the bulk minimum being the zero temperature vacuum of theory, and on only including the leading-order thermal corrections in eq. 27. In the next subsections, we will relax those assumptions.
@@ 4.2.3 Axial or planar minimum as today’s vacuum
In this section we drop our assumption that the zero-temperature vacuum is the bulk extremum, and focus instead on the possibility of an axial or planar minimum today; this corresponds to respectively none or only one of the additional singlets having a vev at zero temperature. We do not attempt a full analysis of the parameter space, but we limit the discussion to point out the conditions that allow for a FOPT. For the sake of readability, we will denote today’s vacuum by simply ‘vacuum’ in this section.
@@
We take $\mathbf{x}_{1}$ as the vacuum. This scenario is not favorable for multi-Higgs boson production enhancement, but we will see that it accommodates a FOPT. Let us consider the transition
@@
With negative masses the planar extremum $\mathbf{x}_{12}$ also exists, but is a saddle point in the $(x_{1},x_{2})$ plane. It follows that there can be a FOPT between the two axial minima through the barrier formed by the potential near $\mathbf{x}_{12}$.
@@
Consider now that the planar extremum $\mathbf{x}_{12}$ is the vacuum, which requires $C_{12}>0$. From the collider point of view, this represents mixing between the Higgs field and only one added scalar, leading to a phenomenology similar to the xSM and a possible enhancement of di-Higgs boson production. For the phase transition, however, there are two distinct possibilities for a FOPT that are unique to the TRSM and have no equivalent in the xSM.
@@
When $\mathbf{x}_{3}$ is a minimum, the other planar extrema $\mathbf{x}_{13},\mathbf{x}_{23}$ do not exist yet. The bulk extremum eq. 34 is however real since all masses $m_{i}^{2}$ are negative. $C_{12}>0$ and $D<0$ imply that the first two eigenvalues of the scaled Hessian in eq. 37 are positive, while the third is negative, meaning that the bulk extremum is a saddle point. We conclude that there can be a FOPT between $\mathbf{x}_{3}$ and $\mathbf{x}_{12}$ with the barrier formed by the potential near the bulk saddle. An illustration of this case is on the right-hand panel of fig. 3. As mentioned before, this case is particularly interesting for LHC phenomenology as a low-temperature vacuum where the Higgs field and one other field have a non-zero vev can lead to an observed enhancement of di-Higgs boson production.
@@
For the bulk vacuum, we saw that the condition eq. 46 prohibits both planar extrema to be a minimum. However, for the planar vacuum, we can again avoid this by taking the determinant eq. 51 negative. For the same parameters as above, the bulk extremum is a saddle which can create a barrier between the two minima at $\mathbf{x}_{12}$ and $\mathbf{x}_{13}$, allowing for a FOPT between them.

### v2

@@ 1 Introduction
“Electroweak baryogenesis” Kuzmin:1985mm ; Shaposhnikov:1986jp ; Shaposhnikov:1987tw ; Cohen:1990py ; Cohen:1993nk is a promising scenario that can generate this asymmetry. It demands the electroweak phase transition to be first order, providing the necessary departure from thermodynamic equilibrium (one of the Sakharov conditions Sakharov:1967dj ). Within the SM, the occurrence of a first-order phase transition (FOPT) requires the mass of the Higgs boson to lie below $\sim 70$ GeV Bochkarev:1987wf ; Kajantie:1995kf ; Kajantie:1996qd ; Csikor:1998eu , which is evidently inconsistent with the experimental observation of the Higgs boson mass of $\sim 125$ GeV. Extended scalar sectors can accommodate an FOPT that is, moreover, sufficiently strong to preserve any baryon asymmetry produced. We note further that the gravitational wave signal from an FOPT at the electroweak scale is in the right frequency range to be potentially observable with the LISA gravitational wave observatory Grojean:2006bp ; Caprini:2018mtu ; Caprini:2019egz ; LISACosmologyWorkingGroup:2022jok .
The SM scalar sector can be extended by an arbitrary number of scalars Chung:2012vg ; Tenkanen:2022tly ; Zuk:2022qwx , but, arguably, the most minimal options are to add a complex or real singlet scalar field Espinosa:1993bs ; Benson:1993qx ; Vergara:1996ub ; Espinosa:2008kw ; Barger:2007im ; Espinosa:2011ax ; Cline:2012hg ; Curtin:2014jma ; Katz:2014bha ; Kanemura:2015fra ; Kurup:2017dzf ; Carena:2019une ; Papaefstathiou:2020iag ; Papaefstathiou:2021glr ; Chen:2022vac ; Azatov:2022tii , two real singlet fields Robens:2019kga ; Papaefstathiou:2020lyp ; Robens:2021zvr ; Robens:2022lkn ; Robens:2022hvu ; Robens:2023pax ; Robens:2023oyz , or additional doublet fields Land:1992sm ; Cline:1996mga ; Turok:1991uc ; Dorsch:2013wja ; Dorsch:2014qja ; Basler:2016obg ; Andersen:2017ika ; Dorsch:2017nza ; Bernon:2017jgv ; Low:2020iua . The additional fields could induce a rich set of phenomena at colliders. In particular, if they mix with the SM-like Higgs boson, their corresponding physical states could be produced singly and decay into fermions or gauge bosons, yielding a signal in multi-lepton or jet searches for heavy resonances.
@@
Triple Higgs boson production is the first attainable process that allows to directly probe the quartic Higgs self-coupling in the scalar potential. Its observation would provide a clear signal for an extended scalar sector, as well as hints on its intricate structure, and possible rôle in electroweak baryogenesis. In fact, one might expect that there exists a correlation between enhanced multi-Higgs boson production and a strong FOPT necessary for baryogenesis, as they both require sufficiently large scalar couplings. Indeed, an FOPT can only arise if the changes to the SM scalar potential are significant, and previous studies of the electroweak phase transition in, e.g. the xSM, shows a clear preference for large couplings Kotwal:2016tex ; Alves:2018oct ; Papaefstathiou:2021glr . This intriguing observation constitutes the main motivation for the present article.
@@
We then study the possibility of an FOPT both in (1) the scenario with the phase transition induced by the temperature-dependence of the leading-order thermal mass corrections, as well as (2) a transition with a radiatively-generated barrier. Surprisingly, we find that the parameter space for large triple Higgs production boson and for an FOPT do not overlap. This result is intricately linked to the $\mathbb{Z}_{2}$-symmetric nature of the set-up and the choice of the vacuum today. Doubly-resonant triple Higgs boson production requires both additional scalars to mix with the Higgs field, and thus both to have a non-zero vacuum expectation value (vev) at zero temperature. If we give up on this requirement, and allow for one or both singlets to have zero vev, an FOPT becomes possible. Like in the xSM, di-Higgs boson production enhancement is possible if only one added scalar has a non-zero vev at low temperature. Unlike the xSM however, the TRSM introduces the possibility of an FOPT resonant di-Higgs boson production.
The present article is organized as follows: in section 2 we discuss the scalar sector model of the $\mathbb{Z}_{2}$-symmetric two real singlet extension of the SM, discuss the conditions for large triple Higgs boson production, and their implications for the perturbativety bound. Section 3 contains our exploration of the parameter space of the model, including a list of updated benchmark points relevant to triple Higgs boson production. In section 4, we review the thermal history of the model, and analyze the possibilities for an FOPT both in the single- and multi-step scenario. We conclude in section 5. Several useful relations for the model are deferred to appendix A, such as the constraints on the boundedness of the potential, the connection between the mass and flavor basis, and the renormalization group equations (RGEs).
@@ 2.1 The two real singlet model
We point out that if one of the singlets has zero vev $v_{i}=0$ for $i=2,3$, the corresponding field $\phi_{i}$ does not mix with the other fields. This scenario does not offer novel collider phenomenology compared to the xSM. However, as we will see in section 4, it does have an interesting effect on cosmology, as it allows for an FOPT.
@@ 3 Updated scan for benchmark points
Figure 2: Scatter plot of the values of $M_{2}$ and $M_{3}$ for the 140 points with triple Higgs boson production cross section over 100 times the SM value. The black solid line denotes the region where double resonant production is kinematically viable, i.e. the boundary $M_{3}=M_{2}+M_{1}$.
@@ 4 Thermal history of the TRSM
It is well-known that loop corrections from particles in the plasma at finite temperature change the shape of the scalar field potential. As a result, the zero-temperature vacuum, in which all three scalar fields have a vev in our model, is not the vacuum state in the early universe. To end up in today’s vacuum, one or more PT(s) had to occur. These PT(s) can be of different types. In an FOPT the vev of one or several fields makes a discontinuous jump, due to a barrier in the potential. Instead, when the fields move continuously from the high-temperature phase to the zero-temperature phase, this is a second-order or cross-over PT. The electroweak PT in the SM is a cross-over Kajantie:1995kf ; Kajantie:1996mn ; Kajantie:1996qd ; Gurtler:1997hr ; Csikor:1998eu ; Aoki:1999fi , but in many BSM models the transition can be of a first-order type, see e.g. Caprini:2019egz and references therein for examples.
An FOPT is phenomenologically very interesting for two reasons: it can source a gravitational wave signal possibly observable by LISA, and it can provide the out-of-equilibrium dynamics necessary for electroweak baryogenesis. Both require the FOPT to be strong; for baryogenesis in particular, the SM Higgs vev should obtain a large value in the transition and
@@
It would be interesting to correlate the parameter space of enhanced triple Higgs production in the TRSM (and enhanced collider phenomenology in general) with the parameters for which a (sufficiently strong) FOPT can occur. For this, we study the effective potential at finite temperature using the framework of thermal field theory. As has become clear in recent years, accurate quantitative statements about PTs require careful many-loop computations Croon:2020cgk ; Niemi:2021qvp ; Gould:2023jbz ; Kierkla:2023von . However, to determine if an FOPT can occur in the TRSM, as a first approximation we can rely on much simpler expressions for the finite-temperature potential, that can be obtained from standard references, such as Dolan:1973qd ; Weinberg:1974hy , or textbooks Kapusta:2006pm ; Bellac:2011kqa .
@@ 4.1 The finite temperature effective potential
$V_{T,{\rm 1-loop}}(\phi,T)=\frac{T^{4}}{2\pi^{2}}\left[\sum_{\alpha=\Phi_{i},W,Z}n_{\alpha}J_{B}[m_{\alpha}^{2}(\phi_{i})/T^{2}]+n_{t}J_{F}[m_{t}^{2}(\phi_{i})/T^{2}]\right],$ (23)
where the masses $m^{2}_{\alpha}$ follow from the tree-level potential.
@@ 4.2 Leading order phase transitions
Note that for the updated benchmark points discussed in Sec. 3, the global minimum does not always lie at the origin at high temperatures in the perturbative regime. Firstly, the scale where the masses become positive may lie above the perturbativity cut-off. Secondly, depending on the couplings, some of the thermal masses in eq. 26 can be negative, leading to a period of ‘symmetry non-restoration’ Weinberg:1974hy ; Mohapatra:1979vr ; Dvali:1995cc ; Meade:2018saz ; Baldes:2018nel ; Matsedonskyi:2020mlz . In the TRSM at least the Higgs thermal mass is always positive, because of the large top quark Yukawa contribution and the boundedness of the negative couplings eq. 54, and thus the electroweak symmetry always gets restored at high temperature. For concreteness, we will consider the case where all thermal masses are positive at temperatures below the perturbativity cutoff, and track the field dynamics as the universe cools down and all fields obtain a vev. If one or multiple thermal masses are already negative at the perturbativity scale, the analysis still applies, but starting from the point where there are already some non-zero vevs. As we will show, one can never have two minima (and thus a FOPT) at the same time, irrespective of the order in which extrema appear. Our results thus do not depend on the assumption that at high temperature all masses are positive and the symmetry is fully restored.
The PT dynamics in the TRSM mirrors that of the xSM, analyzed in e.g. Espinosa:2007qk ; Profumo:2007wc ; Cline:2009sn ; Espinosa:2011ax ; Curtin:2014jma ; Alves:2018jsw ; Gould:2019qek . In the well-studied two-step scenario Land:1992sm , only the Higgs field has a non-zero vev in the vacuum $(h,s)=(v_{h},0)$. Depending on the parameters, it is possible to have a two-step transition, where the first transition $(0,0)\to(0,s)$ is continuous, but the 2nd step $(0,s)\to(h,0)$ is an FOPT. However, if both fields have a vev in the zero-temperature vacuum $(h,s)=(v_{h},v_{s})$, at leading order all possible transitions are second order/cross over as they are always from a saddle point or maximum to a minimum. This has also been demonstrated in Espinosa:2011ax ; Ghorbani:2020xqv .
@@ 4.2.2 Phase transition dynamics
As the temperature drops further, $m_{1}^{2}$ becomes more negative and the extrema separate continuously from each other. The transition from $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$ is thus a continuous PT from a saddle point to a minimum, not an FOPT.
Finally, we note that there can never be two axial minima at the same time. For $m_{i}^{2},m_{j}^{2}<0$ the $\mathbf{x}_{i},\mathbf{x}_{j}$ extrema both exist. A necessary condition for both axial extrema to be a minimum is that $h_{jj}(\mathbf{x}_{i})>0$ and $h_{ii}(\mathbf{x}_{j})>0$, which can only be satisfied for $c_{ij}>0$, which then implies $C_{ij}<0$, which is at odds with our choice of vacuum eq. 38. Therefore, there cannot be an FOPT between different axial minima.
@@
We have already seen that two axial, two planar, and an axial and planar extrema can never be minima at the same time and that the origin is not a minimum once the other extrema emerge. Now our final analysis of the bulk minimum reveals that once it comes into existence, the planar (but also all other) extrema are not minima. Hence, as depicted in the left panel of fig. 3, all possible transitions are smooth and continuous, and no FOPT arises. This conclusion on the absence of an FOPT is based on the bulk minimum being the zero temperature vacuum of theory, and on only including the leading-order thermal corrections in eq. 27. In the next subsections, we will relax those assumptions.
@@ 4.2.3 Axial or planar minimum as today’s vacuum
In this section we drop our assumption that the zero-temperature vacuum is the bulk extremum, and focus instead on the possibility of an axial or planar minimum today; this corresponds to respectively none or only one of the additional singlets having a vev at zero temperature. We do not attempt a full analysis of the parameter space, but we limit the discussion to point out the conditions that allow for an FOPT. For the sake of readability, we will denote today’s vacuum by simply ‘vacuum’ in this section.
@@
We take $\mathbf{x}_{1}$ as the vacuum. This scenario is not favorable for multi-Higgs boson production enhancement, but we will see that it accommodates an FOPT. Let us consider the transition
@@
With negative masses the planar extremum $\mathbf{x}_{12}$ also exists, but is a saddle point in the $(x_{1},x_{2})$ plane. It follows that there can be an FOPT between the two axial minima through the barrier formed by the potential near $\mathbf{x}_{12}$.
@@
Consider now that the planar extremum $\mathbf{x}_{12}$ is the vacuum, which requires $C_{12}>0$. From the collider point of view, this represents mixing between the Higgs field and only one added scalar, leading to a phenomenology similar to the xSM and a possible enhancement of di-Higgs boson production. For the phase transition, however, there are two distinct possibilities for an FOPT that are unique to the TRSM and have no equivalent in the xSM.
@@
When $\mathbf{x}_{3}$ is a minimum, the other planar extrema $\mathbf{x}_{13},\mathbf{x}_{23}$ do not exist yet. The bulk extremum eq. 34 is however real since all masses $m_{i}^{2}$ are negative. $C_{12}>0$ and $D<0$ imply that the first two eigenvalues of the scaled Hessian in eq. 37 are positive, while the third is negative, meaning that the bulk extremum is a saddle point. We conclude that there can be an FOPT between $\mathbf{x}_{3}$ and $\mathbf{x}_{12}$ with the barrier formed by the potential near the bulk saddle. An illustration of this case is on the right-hand panel of fig. 3. As mentioned before, this case is particularly interesting for LHC phenomenology as a low-temperature vacuum where the Higgs field and one other field have a non-zero vev can lead to an observed enhancement of di-Higgs boson production.
@@
For the bulk vacuum, we saw that the condition eq. 46 prohibits both planar extrema to be a minimum. However, for the planar vacuum, we can again avoid this by taking the determinant eq. 51 negative. For the same parameters as above, the bulk extremum is a saddle which can create a barrier between the two minima at $\mathbf{x}_{12}$ and $\mathbf{x}_{13}$, allowing for an FOPT between them.